A car can be slowed down with a constant acceleration of -11 m/s2.


1
Expert's answer
2021-10-10T15:58:35-0400

"a=-11ms^{-1}\\
u=43m/s\\
v=0m/s\\
v^2=u^2+2as\\
s=\dfrac{43^2}{2\times11}=84.1m"